# Radial amplitude equations for fully localised planar patterns

### v1

@@ 1 Introduction
$\sum_{i\in\mathbb{Z}}J_{i}(r)J_{-i}(r)=1.$
However, the proof of Proposition 2.2 is so straightforward that it would be naïve to assume it has not been proven up until now. One aim of this work is then to promote the knowledge of these identities to the wider community and highlight the practicality of using Bessel functions in the study of fully localised planar patterns.
@@ 2.1 Radial differential operators
$\begin{split}\left(1+\Delta_{n}\right)^{2}u_{n}={}&\left(1+\Delta_{n}\right)^{2}u_{n}+\varepsilon\,\left[4\partial_{r}\left(1+\Delta_{n}\right)\partial_{R}u_{n}\right]\\ &+\varepsilon^{2}\left[2\left(1+\Delta_{n}\right)\hat{\Delta}_{0}u_{n}+2\Delta_{n}\hat{\mathcal{D}}_{1}\hat{\mathcal{D}}_{0}u_{n}+\mathcal{D}_{-1-n}\mathcal{D}_{-n}\hat{\mathcal{D}}_{-1}\hat{\mathcal{D}}_{0}u_{n}+\mathcal{D}_{n-1}\mathcal{D}_{n}\hat{\mathcal{D}}_{-1}\hat{\mathcal{D}}_{0}u_{n}\right]+\mathcal{O}(\varepsilon^{3}),\\ ={}&\left(1+\Delta_{n}\right)^{2}u_{n}+\varepsilon\,\left[4\partial_{r}\left(1+\Delta_{n}\right)\partial_{R}u_{n}\right]+\varepsilon^{2}\left[2\left(1+\Delta_{n}\right)\hat{\Delta}_{0}u_{n}+2\Delta_{n}\partial_{R}^{2}u_{n}\right]\\ &+\varepsilon^{2}\left[\left(\mathcal{D}_{n-1}-\mathcal{D}_{1-n}\right)\mathcal{D}_{n}-\left(\mathcal{D}_{n+1}-\mathcal{D}_{-1-n}\right)\mathcal{D}_{-n}\right]\hat{\mathcal{D}}_{-1}\hat{\mathcal{D}}_{0}u_{n}+\mathcal{O}(\varepsilon^{3}),\\ ={}&\left(1+\Delta_{n}\right)^{2}u_{n}+\varepsilon\,\left[4\partial_{r}\left(1+\Delta_{n}\right)\partial_{R}u_{n}\right]+\varepsilon^{2}\left[2\left(1+\Delta_{n}\right)\hat{\Delta}_{0}u_{n}+2\Delta_{n}\partial_{R}^{2}u_{n}\right]+\mathcal{O}(\varepsilon^{3}).\\ \end{split}$
@@
$\begin{split}\left(1+\Delta_{n}\right)^{2}u_{n}={}&\left(1+\Delta_{n}\right)^{2}u_{n}+\varepsilon\,\left[4\partial_{r}\left(1+\Delta_{n}\right)\partial_{R}u_{n}\right]+\varepsilon^{2}\left[2\left(1+\Delta_{n}\right)\hat{\Delta}_{0}u_{n}+2\Delta_{n}\partial_{R}^{2}u_{n}\right]+\mathcal{O}(\varepsilon^{3}).\\ \end{split}$ (2.1)
@@ 2.2 Convolutional sums of Bessel functions
We begin with the following general result for simple combinations of Bessel functions,
@@
Fix $n\in\mathbb{Z}$, $m\in\mathbb{N}$, $j\in\mathbb{N}_{0}$ with $j\leq\frac{m}{2}$, and $\mathbf{k}:=(k_{1},\dots,k_{m})\in\mathbb{Z}^{m}$, with $|\mathbf{k}|:=k_{1}+\dots+k_{m}$. Then,
$J_{n}((m-2j)r)=\sum_{|\mathbf{k}|=n}\;\prod_{i=j+1}^{m}J_{k_{i}}(r)\prod_{\ell=1}^{j}J_{-k_{\ell}}(r),$ (2.2)
where the convolutional sum is taken over all $\mathbf{k}\in\mathbb{Z}^{m}$.
@@
$\mathrm{e}^{\mathrm{i}r\cos(\theta)}=\sum_{k\in\mathbb{Z}}\mathrm{i}^{k}J_{k}(r)\,\mathrm{e}^{\mathrm{i}k\theta}$
@@
$\frac{\mathrm{i}^{-n}}{2\pi}\int_{0}^{2\pi}\mathrm{e}^{\mathrm{i}(r\cos(\theta)-n\theta)}\,\mathrm{d}\theta=J_{n}(r),$ (2.3)
which is known as the Hansen–Bessel formula. We also note the complex conjugate of the Jacobi–Anger expansion
$\mathrm{e}^{-\mathrm{i}r\cos(\theta)}=\sum_{k\in\mathbb{Z}}\mathrm{i}^{-k}J_{k}(r)\,\mathrm{e}^{-\mathrm{i}k\theta}=\sum_{k\in\mathbb{Z}}\mathrm{i}^{k}J_{-k}(r)\,\mathrm{e}^{\mathrm{i}k\theta}$
and consider the product
$\mathrm{e}^{\mathrm{i}(m-2j)r\cos(\theta)}=\left(\prod_{i=j+1}^{m}\sum_{k_{i}\in\mathbb{Z}}\mathrm{i}^{k_{i}}J_{k_{i}}(r)\mathrm{e}^{\mathrm{i}k_{i}\theta}\right)\left(\prod_{\ell=1}^{j}\sum_{k_{\ell}\in\mathbb{Z}}\mathrm{i}^{k_{\ell}}J_{-k_{\ell}}(r)\mathrm{e}^{\mathrm{i}k_{\ell}\theta}\right).$
Collecting the summations, we obtain
$\mathrm{e}^{\mathrm{i}(m-2j)r\cos(\theta)}=\sum_{\mathbf{k}\in\mathbb{Z}^{m}}\mathrm{i}^{|\mathbf{k}|}\mathrm{e}^{\mathrm{i}|\mathbf{k}|\theta}\prod_{i=j+1}^{m}J_{k_{i}}(r)\,\prod_{\ell=1}^{j}J_{-k_{\ell}}(r).$
and so, projecting onto the $n$-th angular Fourier mode, we obtain
$\begin{split}\frac{\mathrm{i}^{-n}}{2\pi}\int_{0}^{2\pi}\mathrm{e}^{\mathrm{i}((m-2j)r\cos(\theta)-n\theta)}\,\mathrm{d}\theta={}&\sum_{\mathbf{k}\in\mathbb{Z}^{m}}\frac{\mathrm{i}^{|\mathbf{k}|-n}}{2\pi}\int_{0}^{2\pi}\mathrm{e}^{\mathrm{i}(|\mathbf{k}|-n)\theta}\,\mathrm{d}\theta\,\prod_{i=j+1}^{m}J_{k_{i}}(r)\,\prod_{\ell=1}^{j}J_{-k_{\ell}}(r),\\ ={}&\sum_{|\mathbf{k}|=n}\;\prod_{i=j+1}^{m}J_{k_{i}}(r)\,\prod_{\ell=1}^{j}J_{-k_{\ell}}(r).\end{split}$
Finally, we note that the left hand side is a particular case of the Hansen–Bessel formula (2.3), and so
$\begin{split}J_{n}((m-2j)r)={}&\sum_{|\mathbf{k}|=n}\;\prod_{i=j+1}^{m}J_{k_{i}}(r)\,\prod_{\ell=1}^{j}J_{-k_{\ell}}(r).\end{split}$
@@
which we will use presently. Clearly each summation is symmetric over $i,j,k\in\mathbb{Z}$, and so one can also permute the summation indices to obtain additional formulae.
@@
Other more complicated patterns can be considered beyond the four that we concentrate on in this paper; for instance one could choose squares or other cases of rotated hexagons considered in [11], including other quasipattern or superlattice structures.
@@ 3 Radial amplitude equation for the Swift–Hohenberg equation
The rest of this section is dedicated to solving (3) by applying the identities introduced in §2.2. We choose particular radial profiles that correspond to the patterns $u_{S},u_{H},u_{R},u_{Q}$ defined in (2.5), (2.6), (2.8) and (2.10), respectively.
@@ 3.1 Fully localised stripes
We restrict to the coefficients of the prime frequency $J_{n}(r)$, where we recover the standard cubic Ginzburg–Landau amplitude equation
@@
The assumption that $A$ is axisymmetric induces a compatibility condition at $R=0$ of the form $\partial_{R}A(0)=0$. By reversibility of the time-independent case of (3.4), any steady localised solution of (3.4) must thus form a homoclinic orbit. Then, for $\nu\leq\sqrt{\frac{27}{38}}$ there are no steady localised solutions of (3.4), and for $\nu>\sqrt{\frac{27}{38}}$ there is a steady localised solution
@@
$\begin{split}\tilde{u}(r,\theta)={}&\varepsilon A(\varepsilon r)\sum_{n\in\mathbb{Z}}\left(J_{n}(r)+J_{-n}(r)\right)\mathrm{e}^{\mathrm{i}n\theta}+\mathcal{O}(\varepsilon^{2}),\\ ={}&2\varepsilon A(\varepsilon r)\sum_{n\in\mathbb{Z}}J_{|2n|}(r)\cos(2n\theta)+\mathcal{O}(\varepsilon^{2}),\\ ={}&2\varepsilon A(\varepsilon r)\cos(r\sin(\theta))+\mathcal{O}(\varepsilon^{2}).\\ \end{split}$
Converting into Cartesian coordinates, we have found a fully localised solution to (1.1) of the form

### v2

@@ 1 Introduction
We note that the use of Bessel functions $J_{n}(r)$ for $n\in\mathbb{Z}$ here is not a particular choice, but rather is enforced by the form of the linear operator $(1+\Delta_{n})^{2}$ of (1.3). In weakly nonlinear analysis, one wants to decompose the solution onto the eigenfunctions of the time-independent linear operator, so that one eigenfunction lies in the kernel of the operator and the remaining elements lie in the range of the operator. This is so that higher order correction terms can be used to remove nonlinear resonances of lower order terms. For example, in one spatial dimension the complex exponential functions $\mathrm{e}^{\mathrm{i}\lambda x}$ act as eigenfunctions of the linear operator $(1+\partial_{x}^{2})^{2}$ and, in particular $(1+\partial_{x}^{2})^{2}\mathrm{e}^{\mathrm{i}x}=0$. Hence, in the one dimensional weakly nonlinear analysis, leading order solutions depend on $\mathrm{e}^{\mathrm{i}x}$ and higher order corrections are decomposed onto the eigenfunctions $\mathrm{e}^{\mathrm{i}\lambda x}$. Returning to the radial linear operator in (1.3), we note that the equation
$(1+\Delta_{n})u_{n}=\left(\partial_{r}^{2}+\frac{1}{r}\partial_{r}+\left(1-\frac{n^{2}}{r^{2}}\right)\right)u_{n}=0$ (1.4)
is Bessel’s equation, and $J_{n}(\lambda r)$ is an eigenfunction of $(1+\Delta_{n})^{2}$ with eigenvalue $(1-\lambda^{2})^{2}$. Hence, the eigenfunction $J_{n}(r)$ lies in the kernel of $(1+\Delta_{n})^{2}$ for each $n\in\mathbb{Z}$, and so we require our leading order solution to depend on $J_{n}(r)$ and higher order corrections to be decomposed onto the eigenfunctions $J_{n}(\lambda r)$.
@@
$\sum_{i\in\mathbb{Z}}J_{i}(r)J_{i}(r)=1.$
However, the derivation of each identity in §2.2 is so straightforward that it would be naïve to assume it has not been proven up until now. One aim of this work is then to promote the knowledge of these identities to the wider community and highlight the practicality of using Bessel functions in the study of fully localised planar patterns.
@@ 2.1 Radial differential operators
$\begin{split}\left(1+\Delta_{n}\right)^{2}u_{n}={}&\left(1+\Delta_{n}\right)^{2}u_{n}+\varepsilon\,\left[4\partial_{r}\left(1+\Delta_{n}\right)\partial_{R}u_{n}\right]\\ &+\varepsilon^{2}\left[2\left(1+\Delta_{n}\right)\hat{\Delta}_{0}u_{n}+2\Delta_{n}\hat{\mathcal{D}}_{1}\hat{\mathcal{D}}_{0}u_{n}+\mathcal{D}_{-1-n}\mathcal{D}_{-n}\hat{\mathcal{D}}_{-1}\hat{\mathcal{D}}_{0}u_{n}+\mathcal{D}_{n-1}\mathcal{D}_{n}\hat{\mathcal{D}}_{-1}\hat{\mathcal{D}}_{0}u_{n}\right]+\mathcal{O}(\varepsilon^{3}),\\ ={}&\left(1+\Delta_{n}\right)^{2}u_{n}+\varepsilon\,\left[4\partial_{r}\left(1+\Delta_{n}\right)\partial_{R}u_{n}\right]+\varepsilon^{2}\left[2\left(1+\Delta_{n}\right)\hat{\Delta}_{0}u_{n}+4\Delta_{n}\partial_{R}^{2}u_{n}\right]\\ &+\varepsilon^{2}\left[\left(\mathcal{D}_{n-1}-\mathcal{D}_{1-n}\right)\mathcal{D}_{n}-\left(\mathcal{D}_{n+1}-\mathcal{D}_{-1-n}\right)\mathcal{D}_{-n}\right]\hat{\mathcal{D}}_{-1}\hat{\mathcal{D}}_{0}u_{n}+\mathcal{O}(\varepsilon^{3}),\\ ={}&\left(1+\Delta_{n}\right)^{2}u_{n}+\varepsilon\,\left[4\partial_{r}\left(1+\Delta_{n}\right)\partial_{R}u_{n}\right]+\varepsilon^{2}\left[2\left(1+\Delta_{n}\right)\hat{\Delta}_{0}u_{n}+4\Delta_{n}\partial_{R}^{2}u_{n}\right]+\mathcal{O}(\varepsilon^{3}).\\ \end{split}$
@@
$\begin{split}\left(1+\Delta_{n}\right)^{2}u_{n}={}&\left(1+\Delta_{n}\right)^{2}u_{n}+\varepsilon\,\left[4\partial_{r}\left(1+\Delta_{n}\right)\partial_{R}u_{n}\right]+\varepsilon^{2}\left[2\left(1+\Delta_{n}\right)\hat{\Delta}_{0}u_{n}+4\Delta_{n}\partial_{R}^{2}u_{n}\right]+\mathcal{O}(\varepsilon^{3}).\\ \end{split}$ (2.1)
@@ 2.2 Convolutional sums of Bessel functions
We begin with the following general result for combinations of Bessel functions,
@@
Fix $n\in\mathbb{Z}$, $m\in\mathbb{N}$, and $\mathbf{k}:=(k_{1},k_{2},\dots,k_{m})\in\mathbb{Z}^{m}$, with $|\mathbf{k}|:=k_{1}+k_{2}+\dots+k_{m}$. Then,
$\sum_{|\mathbf{k}|=n}\,\prod_{i=1}^{m}\mathrm{e}^{\mathrm{i}k_{i}\,\mathrm{arg}(\mathbf{x}_{i})}\,J_{k_{i}}(|\mathbf{x}_{i}|)=J_{n}\left(\left|\sum_{i=1}^{m}\mathbf{x}_{i}\right|\right)\,\mathrm{e}^{\mathrm{i}n\,\mathrm{arg}\left(\sum_{i=1}^{m}\mathbf{x}_{i}\right)}$
holds for all $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{m}\in\mathbb{R}^{2}$, where the convolutional sum is taken over all $\mathbf{k}\in\mathbb{Z}^{m}$, $|\mathbf{x}|$ denotes the modulus of $\mathbf{x}$, and $\arg(\mathbf{x})$ denotes the ‘argument’ of $\mathbf{x}$, i.e. the anticlockwise angle from the positive $x$-axis to the vector $\mathbf{x}$.
@@
$\mathrm{e}^{\mathrm{i}r\cos(\theta)}=\sum_{k\in\mathbb{Z}}J_{k}(r)\,\mathrm{e}^{\mathrm{i}k(\theta+\frac{\pi}{2})}$ (2.2)
@@
$\frac{1}{2\pi}\int_{0}^{2\pi}\mathrm{e}^{\mathrm{i}(r\cos(\theta)-n(\theta+\frac{\pi}{2}))}\,\mathrm{d}\theta=J_{n}(r),$ (2.3)
which is known as the Hansen–Bessel formula. For $\mathbf{x},\mathbf{y}\in\mathbb{R}^{2}$ it follows from (2.2) that
$\mathrm{e}^{\mathrm{i}\mathbf{y}\cdot\mathbf{x}}=\mathrm{e}^{\mathrm{i}|\mathbf{y}||\mathbf{x}|\cos(\mathrm{arg}(\mathbf{x})-\mathrm{arg}(\mathbf{y}))}=\sum_{k\in\mathbb{Z}}J_{k}(|\mathbf{y}|\,|\mathbf{x}|)\,\mathrm{e}^{\mathrm{i}k\,(\mathrm{arg}(\mathbf{x})-\mathrm{arg}(\mathbf{y})+\frac{\pi}{2})}.$ (2.4)
Setting $\mathbf{y}(\varphi)=(\cos(\varphi+\frac{\pi}{2}),\sin(\varphi+\frac{\pi}{2}))^{T}$ for $\varphi\in[0,2\pi)$, we obtain
$\frac{1}{2\pi}\int_{0}^{2\pi}\mathrm{e}^{\mathrm{i}\left(n\varphi+\mathbf{y}(\varphi)\cdot\mathbf{x}\right)}\,\mathrm{d}\varphi=\sum_{k\in\mathbb{Z}}\left(\frac{1}{2\pi}\int_{0}^{2\pi}\mathrm{e}^{\mathrm{i}(n-k)\varphi}\,\mathrm{d}\varphi\right)J_{k}(|\mathbf{x}|)\mathrm{e}^{\mathrm{i}k\,\mathrm{arg}(\mathbf{x})}=J_{n}(|\mathbf{x}|)\,\mathrm{e}^{\mathrm{i}n\,\mathrm{arg}(\mathbf{x})}$
and we can thus derive the generalised Hansen–Bessel formula
$\frac{1}{2\pi}\int_{0}^{2\pi}\mathrm{e}^{\mathrm{i}n\varphi}\,\mathrm{e}^{\mathrm{i}(\cos(\varphi+\frac{\pi}{2}),\sin(\varphi+\frac{\pi}{2}))^{T}\cdot\mathbf{x}}\,\mathrm{d}\varphi=J_{n}(|\mathbf{x}|)\,\mathrm{e}^{\mathrm{i}n\,\mathrm{arg}(\mathbf{x})}.$ (2.5)
Using (2.4) and (2.5), we obtain
$\begin{split}J_{n}\left(\left|\sum_{i=1}^{m}\mathbf{x}_{i}\right|\right)\,\mathrm{e}^{\mathrm{i}n\,\mathrm{arg}\left(\sum_{i=1}^{m}\mathbf{x}_{i}\right)}={}&\frac{1}{2\pi}\int_{0}^{2\pi}\mathrm{e}^{\mathrm{i}n\varphi}\mathrm{e}^{\mathrm{i}(\cos(\varphi+\frac{\pi}{2}),\sin(\varphi+\frac{\pi}{2}))^{T}\cdot\sum_{i=1}^{m}\mathbf{x}_{i}}\,\mathrm{d}\varphi\\ ={}&\frac{1}{2\pi}\int_{0}^{2\pi}\mathrm{e}^{\mathrm{i}n\varphi}\,\prod_{i=1}^{m}\mathrm{e}^{\mathrm{i}(\cos(\varphi+\frac{\pi}{2}),\sin(\varphi+\frac{\pi}{2}))\cdot\mathbf{x}_{i}}\,\mathrm{d}\varphi\\ ={}&\frac{1}{2\pi}\int_{0}^{2\pi}\mathrm{e}^{\mathrm{i}n\varphi}\,\prod_{i=1}^{m}\sum_{\mathbf{k}\in\mathbb{Z}^{m}}J_{k_{i}}(|\mathbf{x}|)\,\mathrm{e}^{\mathrm{i}k_{i}\,(\mathrm{arg}(\mathbf{x}_{i})-\varphi)}\,\mathrm{d}\varphi\\ ={}&\sum_{\mathbf{k}\in\mathbb{Z}^{m}}\left(\frac{1}{2\pi}\int_{0}^{2\pi}\mathrm{e}^{\mathrm{i}\left(n-\sum_{i=1}^{m}k_{i}\right)\varphi}\,\mathrm{d}\varphi\right)\,\prod_{i=1}^{m}J_{k_{i}}(|\mathbf{x}|)\,\mathrm{e}^{\mathrm{i}k_{i}\,\mathrm{arg}(\mathbf{x}_{i})}\\ ={}&\sum_{|\mathbf{k}|=n}\,\prod_{i=1}^{m}J_{k_{i}}(|\mathbf{x}|)\,\mathrm{e}^{\mathrm{i}k_{i}\,\mathrm{arg}(\mathbf{x}_{i})}.\\ \end{split}$
Remark 2.3.
The convolutional identity presented in Proposition 2.2 is a generalisation of Graf’s addition theorem (which can be found in Watson [27, (2-3) in §1̇1.3]), written in vector notation as
$\sum_{i+j=n}\mathrm{e}^{\mathrm{i}\left(i\,\mathrm{arg}(\mathbf{x})+j\,\mathrm{arg}(\mathbf{y})\right)}J_{i}(|\mathbf{x}|)J_{\pm j}(|\mathbf{y}|)={}$ $J_{n}(|\mathbf{x}\pm\mathbf{y}|)\,\mathrm{e}^{\mathrm{i}n\,\mathrm{arg}(\mathbf{x}\pm\mathbf{y})},$ (2.6)
for any $\mathbf{x},\mathbf{y}\in\mathbb{R}^{2}$.
In particular, fixing $j\in[1,m]$ and taking $\mathbf{x}_{i}=(-r,0)^{T}$ for $i\in[1,j]$ and $\mathbf{x}_{i}=(r,0)^{T}$ for $i\in[j+1,m]$, we obtain
$J_{n}((m-2j)r)=\sum_{|\mathbf{k}|=n}\;\prod_{i=j+1}^{m}J_{k_{i}}(r)\prod_{\ell=1}^{j}J_{-k_{\ell}}(r).$ (2.7)
@@
which we will use presently. Here, $\delta_{n,0}$ denotes the standard Kronecker delta, which satisfies $\delta_{i,j}=1$ if $i=j$ and $\delta_{i,j}=0$ otherwise. Clearly each summation is symmetric over $i,j,k\in\mathbb{Z}$, and so one can also permute the summation indices to obtain additional formulae.
@@
Other more complicated patterns can be considered beyond the four that we concentrate on in this paper; for instance one could choose squares or other cases of rotated hexagons considered in [11], including other quasipattern or superlattice structures. In the following remark we highlight an extension of Proposition 2.7 to arbitrary rotations of two hexagon lattices.
Remark 2.8.
Consider the pattern $u_{\alpha}(r,\theta)=\frac{1}{2}\left(u_{H}(r,\theta+\alpha)+u_{H}(r,\theta-\alpha)\right)$ for some $\alpha\in(0,\frac{\pi}{6})$. Then,
$u_{\alpha}(r,\theta)=\sum_{k\in\mathbb{Z}}\mathrm{i}^{3k}c_{k}J_{3k}(r)\,\mathrm{e}^{\mathrm{i}3k\theta}$
where $c_{k}:=\cos(3k\alpha)$, and (2.15) is satisfied by replacing any $b_{n}$ terms by $c_{n}$ and changing the irrational frequencies. In particular, the irrational frequencies $\omega_{1},\dots,\omega_{n}$ within each arbitrary function $f_{i}$ in (2.15) all depend on $\alpha$ but do not equal 1 for any values of $\alpha\in(0,\frac{\pi}{6})$. Hence, an amplitude equation derived for $u_{Q}$ is also an amplitude equation for $u_{\alpha}$ after replacing any $b_{n}$ terms by $c_{n}$ for any $\alpha\in(0,\frac{\pi}{6})$.
@@ 3 Radial amplitude equation for the Swift–Hohenberg equation
The rest of this section is dedicated to solving (3.1) by applying the identities introduced in §2.2. We choose particular radial profiles that correspond to the patterns $u_{S},u_{H},u_{R},u_{Q}$ defined in (2.9), (2.10), (2.12) and (2.14), respectively.
@@ 3.1 Fully localised stripes
We require the $A$-dependent terms to lie in the range of the operator $(1+\Delta_{n})^{2}$, and so the coefficients of the prime frequency $J_{\pm n}(r)$ must be identically zero. Hence, we obtain a solvability condition for $A$ in the form of the standard cubic Ginzburg–Landau amplitude equation
@@
In order to find a steady localised solution of (3.4), we restrict to the invariant subspace $A\in\mathbb{R}$. The assumption that $A$ is axisymmetric induces a compatibility condition at $R=0$ of the form $\partial_{R}A(0)=0$. By reversibility of the time-independent case of (3.4), any steady localised solution of (3.4) must thus form a homoclinic orbit. Then, for $\nu\leq\sqrt{\frac{27}{38}}$ there are no real steady localised solutions of (3.4), and for $\nu>\sqrt{\frac{27}{38}}$ there is a real steady localised solution
@@
$\begin{split}\tilde{u}(r,\theta)={}&\varepsilon A(\varepsilon r)\sum_{n\in\mathbb{Z}}\left(J_{n}(r)+J_{-n}(r)\right)\mathrm{e}^{\mathrm{i}n\theta}+\mathcal{O}(\varepsilon^{2}),\\ ={}&2\varepsilon A(\varepsilon r)\sum_{n\in\mathbb{Z}}J_{|2n|}(r)\cos(2n\theta)+\mathcal{O}(\varepsilon^{2}),\\ ={}&2\varepsilon A(\varepsilon r)\cos(r\sin(\theta))+\mathcal{O}(\varepsilon^{2}),\\ \end{split}$
where we have used the real-valued Jacobi–Anger expansion
$\cos(r\sin(\theta))=\sum_{n\in\mathbb{Z}}J_{|2n|}(r)\,\cos(2n\theta)$
and the fact that $J_{-n}(r)=(-1)^{n}J_{n}(r)$ for any $n\in\mathbb{Z}$. Converting into Cartesian coordinates, we have found a fully localised solution to (1.1) of the form